OpenStudy (anonymous):

Prove that the left side equals the right side. 1 + sec^2x sin^2x = sec^2x

OpenStudy (souvik):

is it sin2x or sin^2x?

OpenStudy (souvik):

1+sin^2x/cos^2x =1+tan^2x =sec ^2x

OpenStudy (anonymous):

OK wow that's good. HOw did you get the 1/cos^2? Is that the reciprocal of sec^2?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (souvik):

yes...secx=1/cosx so sec^2x=1/cos^2x

OpenStudy (anonymous):

Oh i'm sorry i chose the wrong best answer. Thanks anyways jdoe